# Characterization of Nanostructural Imperfections in Superconducting Quantum Circuits

### v1

@@ Abstract
Decoherence in superconducting quantum circuits, caused by loss mechanisms like material imperfections and two-level system (TLS) defects, remains a major obstacle to improving the performance of quantum devices. In this work, we present atomic-level characterization of cross-sections of a Josephson junction and a spiral resonator to assess the quality of critical interfaces. Employing scanning transmission electron microscopy (STEM) combined with energy-dispersive X-ray spectroscopy (EDS) and electron-energy loss spectroscopy (EELS), we identify structural imperfections associated with oxide layer formation and carbon-based contamination, and correlate these imperfections to the pattering and etching steps in the fabrication process and environmental exposure. These results help to understand that TLS imperfections at critical interfaces play a key role in limiting device performance, emphasizing the need for an improved fabrication process.
@@ Introduction
Improvements in the MS interface have been achieved through substrate cleaning via oxide etching before thin-film growth [19], controlled epitaxial film deposition [25], and interface-targeted etching techniques for resonator and qubit structures [26, 27]. However, controlling native oxide formation at MA and SA interfaces remains challenging, as thin films and substrates are typically exposed to ambient conditions. Chemical etching of oxides at SA and MA interfaces can solve part of the problem and significantly improve resonator quality factors [28], however this is limited by the rapid regrowth of oxide layers.
@@
In this work, we present detailed materials characterization of the metal-substrate (MS), metal-air (MA), and substrate-air (SA) interfaces of a superconducting quantum device fabricated at University of Oxford, focusing on oxide layer formation and carbon contamination. Using high-resolution cross-sectional imaging with scanning transmission electron microscopy (STEM) combined with energy-dispersive X-ray spectroscopy (EDS) and electron energy-loss spectroscopy (EELS), we examine a Josephson junction (JJ) and a spiral resonator. The nanostructural imperfections identified are correlated with the pattering and etching steps in the fabrication process and environmental exposure [14, 28]. Our study empathizes on the need of using materials characterization techniques for assessing nano-scale surface roughness and interface oxidation, providing critical insights about nanostructural imperfections in superconducting quantum devices.
@@
The detailed steps of the fabrication process are described [40, 41], with relevant design parameters given in Table 1. Spin-coating a protective photoresist layer on the backside is critical in double-sided fabrication process to protect the wafer and prevent additional contamination. The resonators side is patterned first while the qubit side is covered with photoresist, followed by cleaning the photoresist and spin-coating another protective photoresist layer on the resonators side and pattering the qubits side.
@@
Immediately after water cleaning, the wafer is immediately loaded in Plassys MEB550S2 at ultra-high vacuum (UHV) and is baked up to 200 C for 10 min. After which a layer of 100 nm of aluminum is deposited at rate of 1 nm/s on the substrate through UHV electron-beam evaporation under controlled temperature and low-pressure conditions, with a base pressure down to mbar and an evaporation pressure of around mbar, ensuring high purity and uniformity of the thin film. The deposition rate and substrate temperature are carefully controlled to ensure smooth thin-film growth for optimal grain structure.
@@
Electron-Beam Lithography and Non-scale Josephson Junctions
@@
After EBL patterning, the wafer is loaded into the Plassys MEB550S2. Prior to deposition, an argon (Ar) ion milling is performed for 1 min (voltage 400 V, acceleration voltage 90 V/s, current 15 mA) to remove any residual contaminants and native oxides from the metal and substrate surfaces, ensuring a clean interface for the subsequent aluminum deposition. The first layer of junction is then deposited at an angle of from normal incidence, depositing 60 nm of Al at a rate of 0.5 nm/s. Due to the deposition angle, the effective thickness of the deposited film is approximately 27-30 nm. Following the first deposition, an in situ controlled static oxidation inside Plassys is performed, typically for 5-10 min at an oxygen pressure of 5-10 mbar, depending on the target junction resistance. This controlled oxidation forms the thin insulating barrier of aluminum oxide essential for the tunnel junction. After pumping back down to UHV conditions, the second layer of aluminum is deposited at normal incidence (), depositing 70 nm of Al at a rate of 0.5 nm/s, effectively completing the Josephson junction structure. Precise control over the oxidation parameters, such as oxygen pressure and exposure time, is critical to achieve the desired tunnel barrier properties and, consequently, the critical current of the junction [42]. Following evaporation, a lift-off process is carried out in a DMSO solution at 80 C for around 2 hrs and immediately followed by thoroughly rinsing with ultrapure deionized water and drying using nitrogen gas.
@@
Cross-sectional lamellae were milled to expose the regions of interest, lifted out, and transferred onto TEM grids at UHV. These lamellae were then carefully thinned using progressively lower ion beam currents, with final polishing at 10 kV and 5 kV to minimize amorphization and enhance imaging quality shown in Fig. 1(c) and Fig. 1(d). The samples were transferred out of the FIB-SEM under vacuum conditions using a transfer holder and placed in a glovebox to prevent oxidation. This approach ensured that the structural and compositional integrity of the samples was maintained prior to materials analysis. By minimizing exposure to ambient conditions, we minimized the possibility of introducing additional surface oxidation or contamination, which could interfere with the desired characterization.
@@
Scanning transmission electron microscopy (STEM) was performed at the electron Physical Science Imaging Centre (ePSIC). The qubit side of the device was analysed with a JEOL ARM300CF aberration-corrected instrument (E02) equipped with an Oxford Instruments XMAX 100 energy-dispersive X-ray spectroscopy (EDS) detector. The resonator side of the device was analysed with a JEOL ARM200CF aberration-corrected instrument (E01) equipped with a Gatan Dual electron-energy loss spectroscopy (EELS) detector and a large solid angle EDX detector (JEOL Centurio). The ARM200CF operated with a convergence semi-angle of 23 mrad, an EELS collection semi-angle of 52 mrad and a probe current of 41 pA. The EELS and EDX data were acquired using the Gatan Microscopy Suite and analysed using Hyperspy software. The plasmon background of the EELS were removed by fitting a power law to the region before the oxygen edge in the 450-500 eV range. The O K-edge, Al K-edge and Si K-edge components in the spectra were fitted using open source generalised oscillator strengths [43] to compute the double differential cross sections for inelastic electron scattering by atoms. The EDX intensities were obtained using integration windows of width 3 eV, carefully selecting suitable background windows to avoid peak overlaps. EELS and EDX spectra from regions of interest of different size have been normalised by summing the intensity over the region and dividing by its area. Linescans have been created from spectrum images by summing data in a series of 1 nm thick slices with widths indicated on the images.
@@
The superconducting quantum device studied in this work consists of transmon qubits arranged in a lattice within a 3D-integrated coaxial circuit quantum electrodynamics (cQED) architecture [44, 45, 46]. In this section, we present a detailed analysis on two critical components on both sides of the fabricated device: the Josephson junction (qubit side in Fig. 1(a)) and the spiral inductor (resonator side in Fig. 1(b)).
@@
The low carbon contamination in the AlO tunnel barrier (ROI-2) suggests that the controlled oxidation process effectively produces a high purity insulating layer essential for the junction. However, the pronounced oxygen presence at the MA and SA interfaces indicates that surface oxidation is occurring post-fabrication, likely due to the subsequent fabrication process and exposure to ambient conditions. The carbon levels across different ROIs in the left section of the lamella imply that carbon contamination could be widespread, possibly introduced during sample preparation or from residual processing materials. This pervasive carbon and oxygen at air interfaces could contribute to dielectric loss, building a decoherence channel for the qubit. Implementing improved surface cleaning and protective measures during and after the fabrication process could be vital in mitigating these unwanted oxidation and contamination.
In addition, linescans reconstructed from EDX maps across the surface oxide layers of the Josephson junction area are shown in in Fig. 3(b) and (c). Line-1 and Line-2 were analyzed to assess the thickness and composition of the oxide layers in different regions. In Line-1, the surface oxide layer is approximately 10 nm thick. The aluminum content appears to increase with depth into the surface; however, here the layers are not completely flat which may cause some smearing of the features. Additionally, there is a noticeable presence of carbon at the very top of the oxide layer within the platinum capping layer, but less within the underlying device. In contrast, Line-2 reveals an oxide layer of about 7 nm thick exhibiting low levels of carbon contamination.
@@
Next, we discuss materials analysis on the spiral inductor in Fig. 1(b), where the results in Fig. 4 presents a detailed analysis of the middle section of the lamella, a key component in the resonator circuit. In Fig. 4(a), the cross-sectional STEM-EELS maps highlight regions of interest (ROIs) across the interface, marked by boxes. The individual images for (b) oxygen, aluminum, and silicon illustrate the distribution of these elements in the examined region. The corresponding EELS maps in Figure 4(c) clearly show three distinct regions: AlOx surface (ROI-1), Al film (ROI-2), and Si substrate (ROI-3). In addition, linescans reconstructed from EDX maps along different regions are presented in Figures 4(d) and 4(e), corresponding to Line-1, Line-2, and Line-3, respectively. The elemental profiles along Line-1 reveal the presence of a surface oxide layer approximately 10 nm thick, characterized by significant concentrations of aluminum and oxygen. In contrast, the profiles along Line-2 show a much cleaner interface with no relative oxygen signal, suggesting minimal oxidation at this boundary with sharper aluminum and silicon profiles. These findings highlight the variability in interface quality as the differentiation between oxidized and cleaner regions indicate the importance of precise fabrication techniques to control the oxidation process and maintain the integrity of the interface.
@@
These observations indicate that the sidewalls and base of the aluminum thin-film are more prone to oxidation and contamination compared to the top surface. The increased oxygen content from top to side to base suggests that these regions accumulate thicker oxide layers. The higher carbon content at the base and sides, as revealed by EDX, implies the presence of organic impurities. This suggests that photolithography and wet-etching steps which selectively target these regions may leave behind contaminants that preferentially adhere to the sidewalls and base. The cleaner oxide layer on the top surface indicates that this area is less exposed to contaminants or that any impurities are more effectively removed or prevented from adhering during processing. These differences in impurity levels are critical because the thicker, impurity-rich oxide layers at the side and base could introduce additional dielectric loss. Further optimizing the photolithography and etching processes to reduce these imperfections, especially at the sidewalls and base, is essential for enhancing the quality of the processed aluminum thin-film device.
@@
Furthermore, Fig. 6 provides a detailed compositional analysis of the right side of the cross-section. In Fig. 6(a), the EELS maps highlights the distribution materials composition and the EELS maps in Fig. 6(b) and 6(c) display the localized presence of oxygen and aluminum, respectively, confirming the formation of aluminum oxide (AlO) layers. Linescans reconstructed from EELS maps along three different regions are presented in Fig. 6(d), 6(e), and 6(f), corresponding to Line-1, Line-2, and Line-3, respectively. The elemental profiles along Line-1 reveal a surface oxide layer, with increasing aluminum content as depth increases, suggesting a gradient in the thickness of the oxide layer from the surface inward. The profiles along Line-2 indicate thicker oxide layer with an increase in silicon content at depth. This may indicate interdiffusion or oxidation extending into both the metal and substrate. In contrast, the profiles along Line-3 show a much cleaner interface, with no relative oxygen signal and sharp transitions in aluminum and silicon content across the interface, suggesting minimal oxidation and a well-defined boundary between the aluminum thin film and the silicon substrate.
@@
Additionally, in Fig. 7(a), the cross-sectional STEM images highlight the distribution of materials composition at further right of the lamella. The individual elemental maps of the same area confirm that this bulk of the residual layer observed on both sides of the resonator track is predominantly composed of oxygen, rather than aluminum oxide (AlO), silicon oxide (SiO) or a combination of both. This indicates that oxidation is occurring mainly within this much thinner aluminum film rather than at the aluminum-silicon interface. From the linescans reconstructed along two different regions in Fig 7(a) we can see the elemental profiles in which Line-1 (see Fig 7(e)) shows concentrations of oxygen, revealing a surface oxide layer approximately 20 nm thick with increasing silicon content as depth increases. This suggests a substantial oxidation layer on the surface in the aluminum thin film. In contrast, the profile along Line-2, which is further to the right, shows non-uniform but thinner oxide layers with increasing silicon content with depth. This may indicate some degree of interfacial oxidation or diffusion of oxygen into the silicon substrate in this region.

### v2

@@ Abstract
Decoherence in superconducting quantum circuits, caused by loss mechanisms like material imperfections and two-level system (TLS) defects, remains a major obstacle to improving the performance of quantum devices. In this work, we present atomic-level characterization of cross-sections of a Josephson junction and a spiral resonator to assess the quality of critical interfaces. Employing scanning transmission electron microscopy (STEM) combined with energy-dispersive X-ray spectroscopy (EDS) and electron-energy loss spectroscopy (EELS), we identify structural imperfections associated with oxide layer formation and carbon-based contamination, and correlate these imperfections to the patterning and etching steps in the fabrication process and environmental exposure. These results suggest that TLS imperfections at critical interfaces significantly contribute to limiting device performance, emphasizing the need for an improved fabrication process.
@@ Introduction
Improvements in the MS interface have been achieved through substrate cleaning by oxide etching before thin-film growth [19], controlled epitaxial film deposition [25], and interface-targeted etching techniques for resonator and qubit structures [26, 27]. However, controlling native oxide formation at MA and SA interfaces remains challenging, as thin films and substrates are typically exposed to ambient conditions. Chemical etching of oxides at SA and MA interfaces can solve part of the problem and significantly improve resonator quality factors [28], however this is limited by the rapid regrowth of oxide layers.
@@
In this work, we present detailed materials characterization of the metal-substrate (MS), metal-air (MA), and substrate-air (SA) interfaces of a superconducting quantum device fabricated at University of Oxford, focusing on oxide layer formation and carbon contamination. Using high-resolution cross-sectional imaging with scanning transmission electron microscopy (STEM) combined with energy-dispersive X-ray spectroscopy (EDS) and electron energy-loss spectroscopy (EELS), we examine a Josephson junction (JJ) and a spiral resonator. The nanostructural imperfections identified are correlated with the patterning and etching steps in the fabrication process and environmental exposure [14, 28]. Our study emphasizes on the need of using materials characterization techniques for assessing nano-scale surface roughness and interface oxidation, providing critical insights about nanostructural imperfections in superconducting quantum devices.
@@
The superconducting quantum device studied in this work consists of transmon qubits arranged in a lattice within a 3D-integrated coaxial circuit quantum electrodynamics (cQED) architecture [40, 41, 42]. The qubits and resonators are fabricated on opposite sides of the silicon substrate and capacitively coupled through the bulk substrate. The coupling strength is primarily determined by the substrate thickness and the geometry of the capacitive pads of both the qubit (in Fig. 1(a)) and the resonator (in Fig. 1(b)) [40]. The detailed steps of the fabrication process are described [43, 44], with relevant design parameters given in Table 1. Spin-coating a protective photoresist layer on the backside is critical in double-sided fabrication process to protect the wafer and prevent additional contamination. The resonators side is patterned first while the qubit side is covered with photoresist, followed by cleaning the photoresist and spin-coating another protective photoresist layer on the resonators side and patterning the qubits side.
@@
Immediately after water cleaning, the wafer is immediately loaded in Plassys MEB550S2 at ultra-high vacuum (UHV) and is baked up to for 10 min. After which a layer of 100 nm of aluminum is deposited at rate of 1 nm/s on the substrate through UHV electron-beam evaporation under controlled temperature and low-pressure conditions, with a base pressure down to mbar and an evaporation pressure of around mbar, ensuring high purity and uniformity of the thin film. The deposition rate and substrate temperature are carefully controlled to ensure smooth thin-film growth for optimal grain structure.
@@
Electron-Beam Lithography and Nano-scale Josephson Junctions
@@
After EBL patterning, the wafer is loaded into the Plassys MEB550S2. Prior to deposition, an argon (Ar) ion milling is performed for 1 min (voltage 400 V, acceleration voltage 90 V/s, current 15 mA) to remove any residual contaminants and native oxides from the metal and substrate surfaces, ensuring a clean interface for the subsequent aluminum deposition. The first layer of junction is then deposited at an angle of from normal incidence, depositing 60 nm of Al at a rate of 0.5 nm/s. Due to the deposition angle, the effective thickness of the deposited film is approximately 27-30 nm. Following the first deposition, an in situ controlled static oxidation inside Plassys is performed, typically for 5-10 min at an oxygen pressure of 5-10 mbar, depending on the target junction resistance. This controlled oxidation forms the thin insulating barrier of aluminum oxide essential for the tunnel junction. After pumping back down to UHV conditions, the second layer of aluminum is deposited at normal incidence (), depositing 70 nm of Al at a rate of 0.5 nm/s, effectively completing the Josephson junction structure. Precise control over the oxidation parameters, such as oxygen pressure and exposure time, is critical to achieve the desired tunnel barrier properties and, consequently, the critical current of the junction [45]. Following evaporation, a lift-off process is carried out in a DMSO solution at for around 2 hrs and immediately followed by thoroughly rinsing with ultrapure deionized water and drying using nitrogen gas.
@@
Cross-sectional lamellae were milled to expose the regions of interest, lifted out, and transferred onto TEM grids under a high vacuum of approximately .. These lamellae were then carefully thinned using progressively lower ion beam currents, with final polishing at 10 kV and 5 kV to minimize amorphization and enhance imaging quality shown in Fig. 2(c) and Fig. 2(d). The samples were transferred out of the FIB-SEM under vacuum conditions using a transfer holder and placed in a glovebox to prevent oxidation. This approach ensured that the structural and compositional integrity of the samples was maintained prior to materials analysis. By minimizing exposure to ambient conditions, we minimized the possibility of introducing additional surface oxidation or contamination, which could interfere with the desired characterization.
@@
Scanning transmission electron microscopy (STEM) was performed at the electron Physical Science Imaging Centre (ePSIC). The qubit side of the device was analyzed with a JEOL ARM300CF aberration-corrected instrument (E02) equipped with an Oxford Instruments XMAX 100 energy-dispersive X-ray spectroscopy (EDS) detector. The resonator side of the device was analyzed with a JEOL ARM200CF aberration-corrected instrument (E01) equipped with a Gatan Dual electron-energy loss spectroscopy (EELS) detector and a large solid angle EDX detector (JEOL Centurio). The ARM200CF operated with a convergence semi-angle of 23 mrad, an EELS collection semi-angle of 52 mrad and a probe current of 41 pA. The EELS and EDX data were acquired using the Gatan Microscopy Suite and analyzed using Hyperspy software [46]. The plasmon background of the EELS were removed by fitting a power law to the region before the oxygen edge in the 450-500 eV range. The O K-edge, Al K-edge and Si K-edge components in the spectra were fitted using open source generalised oscillator strengths [47] to compute the double differential cross sections for inelastic electron scattering by atoms. The EDX intensities were obtained using integration windows of width 3 eV, carefully selecting suitable background windows to avoid peak overlaps. EELS and EDX spectra from regions of interest of different size have been normalised by summing the intensity over the region and dividing by its area. Linescans have been created from spectrum images by summing data in a series of 1 nm thick slices with widths indicated on the images.
@@
In this section, we present a detailed analysis on two critical components on both sides of the fabricated device: the Josephson junction (qubit side in Fig. 2(a)) and the spiral inductor (resonator side in Fig. 2(b)). The analysis on oxide thickness, elemental distribution, and surface roughness discussed here represent qualitative trends observed across a limited number of cross-sectional lamellae. Given the inherent non-uniformity of surface oxidation, carbon contamination, and roughness across the device, combined with the localized nature of the TEM-based analysis, a detailed quantitative analysis would not be robust and reproducible. For this reasons, we focus this discussion on spatial trends and the relative differences between interfaces, while including limited quantitative analysis through linescans across the interfaces, rather than stressing on absolute quantitative measures.
@@
The low carbon contamination in the AlO tunnel barrier (ROI-2) suggests that the controlled oxidation process effectively produces a high purity insulating layer essential for the junction. However, the pronounced oxygen presence at the MA and SA interfaces indicates that surface oxidation is occurring post-fabrication, likely due to the subsequent fabrication process and exposure to ambient conditions. The carbon levels across different ROIs in the left section of the lamella imply that carbon contamination could be widespread, possibly introduced during sample preparation or from residual processing materials. This pervasive carbon and oxygen at air interfaces could contribute to dielectric loss, building a decoherence channel for the qubit.
Both carbon contamination and oxide formation at interfaces potentially originate from several sources, including residual organic films left after lithography and resist stripping, and handling-related contamination introduced during processing and sample preparation [7, 8]. These residues contribute to the formation of amorphous dielectric layers, which are known to host TLS defects that can couple to both qubits resonators, acting as parasitic loss channel that absorb and dissipate microwave energy, limiting both qubit coherence times and resonator quality factors [28]. Methods to mitigate carbon contamination and TLS loss include optimized resist removal using solvent cleaning combined with in-situ plasma ashing, minimized ambient exposure through vacuum or inert gas transfer between processing steps, and surface passivation immediately after etching to prevent further contamination [34, 35]. Implementing improved surface cleaning and protective measures during and after the fabrication process could be vital in mitigating these unwanted oxidation and contamination.
In addition, linescans reconstructed from EDX maps across the surface oxide layers of the Josephson junction area are shown in Fig. 4(b) and (c). Line-1 and Line-2 were analyzed to assess the thickness and composition of the oxide layers in different regions. In Line-1, the surface oxide layer is approximately 10 nm thick. The aluminum content appears to increase with depth into the surface; however, here the layers are not completely flat which may cause some smearing of the features. Additionally, there is a noticeable presence of carbon at the very top of the oxide layer within the platinum capping layer, but less within the underlying device. In contrast, Line-2 reveals an oxide layer of about 7 nm thick exhibiting low levels of carbon contamination.
@@
Next, we discuss materials analysis on the spiral inductor in Fig. 2(b), where the results in Fig. 5 presents a detailed analysis of the middle section of the lamella, a key component in the resonator circuit. In Fig. 5(a), the cross-sectional STEM-EELS maps highlight regions of interest (ROIs) across the interface, marked by boxes. The individual images for (b) oxygen, aluminum, and silicon illustrate the distribution of these elements in the examined region. Note: the observed silicon signal in the platinum capping layer in (b) is likely an artefact from lamella projection. The corresponding EELS maps in Figure 5(c) clearly show three distinct regions: AlOx surface (ROI-1), Al film (ROI-2), and Si substrate (ROI-3). In addition, linescans reconstructed from EDX maps along different regions are presented in Figures 5(d) and 5(e), corresponding to Line-1, Line-2, and Line-3, respectively. The elemental profiles along Line-1 reveal the presence of a surface oxide layer approximately 10 nm thick, characterized by significant concentrations of aluminum and oxygen. In contrast, the profiles along Line-2 show a much cleaner interface with no relative oxygen signal as also seen in Fig. 5(b), suggesting minimal oxidation at this boundary with sharper aluminum and silicon profiles. This observation of a relatively clean MS interface with minimal oxygen content is consistent with the Josephson junction results shown in Fig. 3(b), where the MS interface was also found to be notably clean across the sample. These findings highlight the variability in interface quality as the differentiation between oxidized and cleaner regions indicate the importance of precise fabrication techniques to control the oxidation process and maintain the integrity of the interface.
@@
These observations indicate that the sidewalls and base of the aluminum thin-film are more prone to oxidation and contamination compared to the top surface. The increased oxygen content from top to side to base suggests that these regions accumulate thicker oxide layers. The higher carbon content at the base and sides, as revealed by EDX, implies the presence of organic impurities. This suggests that photolithography and wet-etching steps which selectively target these regions may leave behind contaminants that preferentially adhere to the sidewalls and base. While the aluminum intensity appears similar across all three regions (see Fig. 6(a)), the oxygen signal increases progressively from the top to the base. This suggests that the AlOx in these areas originates from the same oxidation mechanism during fabrication and ambient exposure, but varies in oxygen level and thickness likely due to spatially non-uniform etching dynamics, particularly at the base.. The cleaner oxide layer on the top surface indicates that this area is less exposed to contaminants or that any impurities are more effectively removed or prevented from adhering during processing. These differences in impurity levels are critical because the thicker, impurity-rich oxide layers at the side and base could introduce additional dielectric loss. Further optimizing the photolithography and etching processes to reduce these imperfections, especially at the sidewalls and base, is essential for enhancing the quality of the processed aluminum thin-film device.
@@
Furthermore, Fig. 7 provides a detailed compositional analysis of the right side of the cross-section. In Fig. 7(a), the EELS maps highlights the distribution materials composition and the EELS maps in Fig. 7(b) and 7(c) display the localized presence of oxygen and aluminum, respectively, confirming the formation of aluminum oxide (AlO) layers. Linescans reconstructed from EELS maps along three different regions are presented in Fig. 7(d), 7(e), and 7(f), corresponding to Line-1, Line-2, and Line-3, respectively. The elemental profiles along Line-1 reveal a surface oxide layer, with increasing aluminum content as depth increases, suggesting a gradient in the thickness of the oxide layer from the surface inward. The profiles along Line-2 indicate thicker oxide layer with an increase in silicon content at depth. This may indicate interdiffusion or oxidation extending into both the metal and substrate. In contrast, the profiles along Line-3 show a much cleaner interface, with no relative oxygen signal and sharp transitions in aluminum and silicon content across the interface, suggesting minimal oxidation and a well-defined boundary between the aluminum thin film and the silicon substrate.
@@
Additionally, in Fig. 8(a), the cross-sectional STEM images highlight the distribution of materials composition at further right of the lamella. The individual elemental maps of the same area confirm that this bulk of the residual layer observed on both sides of the spiral inductor is predominantly composed of oxygen with only a thin presence of aluminum, indicating that it is primarily aluminum oxide (AlOx) rather than silicon oxide (SiOx). This indicates that oxidation is occurring mainly within this much thinner aluminum film rather than at the aluminum-silicon interface. From the linescans reconstructed along two different regions in Fig 8(a) we can see the elemental profiles in which Line-1 (see Fig 8(e)) shows concentrations of oxygen, revealing a surface oxide layer approximately 20 nm thick with increasing silicon content as depth increases. This suggests a substantial oxidation layer on the surface in the aluminum thin film. In contrast, the profile along Line-2, which is further to the right, shows non-uniform but thinner oxide layers with increasing silicon content with depth. This may indicate some degree of interfacial oxidation or diffusion of oxygen into the silicon substrate in this region.
Microwave Characterization
Basic microwave characterization and device parameters of 16 qubits and 16 resonators from the same wafer are shown in Table. 2, using a standard microwave and cryogenic setup described here [40, 41, 42]. Resonators are designed to have distinct frequencies following a ladder design ranging from ~ 8.6 to 10.0 GHz. Each set of 8 qubits are designed with two-distinct, alternating, frequency pattern in a range between 4.8 and 4.9 GHz, with very low weighted-frequency spreads below 0.5 and 1.5 MHz, receptively, for both targeted values. The average anharmonicity is MHz across all qubits with a weighted-frequency spread of 0.3 MHz. The low frequency spreads were achieved without any further post-fabrication process on the junctions, such as junctions annealing [48], but only by fine tuning the junctions fabrication parameters.
Qubits relaxation and coherence times were measured repeatedly over 12 hours, resulting in a total of 400 measurements for each , and that are then averaged for each qubit. The average qubits relaxation times , and dephasing times and were measured using Ramsey and Hahn echo sequences, respectively. Relaxation and coherence times averaged across the device are , and , with weighted-standard deviations.
While a direct quantitative correlation between atomic-scale material imperfections and individual qubit performance is not established here, the inclusion of microwave characterization provides a meaningful benchmark for circuits fabricated using the same process flow analyzed in this study. The coherence times and quality factors reported across the 16 qubits and resonators offer insight into the level of performance achievable in the presence of the types of structural imperfections observed using the presented materials characterization techniques. We emphasize that although localized defects (such as interface oxidation or carbon contamination) may vary in position and extent, the fabrication process is consistent across the device, allowing us to infer a representative range of performance outcomes resulting from these imperfections.
| Parameters | |||||||||
| Qubits | MHz | MHz | MHz | KHz | MHz | ||||
| 9997.4 | 4888.2 | 11.7 | 2.6 | -200.0 | -196.6 | 126 ± 18 | 107 ± 12 | 124 ± 23 | |
| 9386.0 | 4795.6 | 11.8 | 1.3 | -225.0 | -197.2 | 89 ± 13 | 56 ± 15 | 86 ± 12 | |
| 9299.2 | 4807.5 | 6.5 | 1.8 | -200.0 | -196.2 | 61 ± 6 | 44 ± 5 | 102 ± 18 | |
| 8649.5 | 4809.8 | 5.3 | 2.8 | -200.0 | -198.6 | 54 ± 9 | 39 ± 14 | 97 ± 20 | |
| 8755.6 | 4855.3 | 3.1 | 0.8 | -225.0 | -196.4 | 68 ± 10 | 38 ± 4 | 45 ± 10 | |
| 9220.6 | 4824.8 | 6.4 | 0.5 | -225.0 | -194.0 | 63 ± 7 | 49 ± 4 | 68 ± 10 | |
| 9474.2 | 4928.5 | 10.8 | 1.7 | -175.0 | -195.6 | 77 ± 12 | 45 ± 12 | 82 ± 15 | |
| 9908.6 | 4829.5 | 5.8 | 2.0 | -175.0 | -197.2 | 63 ± 8 | 32 ± 7 | 71 ± 8 | |
| 9802.3 | 4963.4 | 7.8 | 4.4 | -250.0 | -195.0 | 24 ± 7 | 24 ± 5 | 33 ± 9 | |
| 9535.4 | 4817.2 | 7.3 | 3.2 | -200.0 | -196.9 | 51 ± 10 | 35 ± 9 | 63 ± 17 | |
| 9112.8 | 4835.4 | 16.5 | 1.4 | -175.0 | -196.1 | 74 ± 7 | 49 ± 3 | 78 ± 13 | |
| 8851.8 | 4777.3 | 11.4 | 0.9 | -250.0 | -196.4 | 92 ± 24 | 57 ± 11 | 76 ± 15 | |
| 8943.2 | 4884.0 | 16.7 | 1.6 | -175.0 | -195.3 | 102 ± 13 | 63 ± 7 | 107 ± 22 | |
| 9025.3 | 4855.2 | 9.3 | 1.4 | -175.0 | -197.0 | 56 ± 12 | 56 ± 10 | 60 ± 15 | |
| 9645.5 | 5040.2 | 7.3 | 1.7 | -225.0 | -196.1 | 55 ± 7 | 58 ± 9 | 65 ± 9 | |
| 9728.7 | 4792.8 | 24.6 | 2.7 | -175.0 | -197.5 | 60 ± 6 | 46 ± 4 | 65 ± 9 | |
| Statistics | |||||||||
| Max | 9997.4 | 5040.2 | 24.6 | 4.4 | -175.0 | -194.0 | 126 | 107 | 124 |
| Min | 8649.5 | 4777.3 | 3.1 | 0.5 | -250.0 | -198.6 | 41 | 32 | 45 |
| (Mean) | 9333.5 | 4856.5 | 10.1 | 1.9 | -203.1 | -196.4 | 71 | 51 | 78 |
| (Std. Dev) | 407.12 | - | 5.3 | 0.96 | 26.3 | 1.1 | 21 | 17 | 20 |
| () | 101.8 | - | 1.3 | 0.2 | 6.6 | 0.3 | 5 | 4 | 5 |